# Extraordinary surface critical behavior induced by a topological dimer phase in a two-dimensional Heisenberg quantum magnet

### v1

@@ Abstract
Using quantum Monte Carlo simulations, we study the dimerized spin-1/2 Heisenberg model on a square lattice, focusing on the paramagnetic dimer phase and its edge states, and associated surface criticality behavior. The model can also be viewed as two-dimensional antiferromagnetically (AF) coupled usual ladders with AF rungs. We demonstrate that the dimer phase of the model, previously considered topologically trivial, is actually a quasi-one-dimensional (Q1D) Haldane phase belonging to a symmetry-protected topological (SPT) state. This is established by measuring generalized strange correlators introduced in this work and further confirmed by showing that the nontrivial edge state on a zigzag cut in the dimer phase is ferromagnetically ordered, resulting from effective ferromagnetic interactions between degenerate spinons generated on each side of the zigzag cut. Furthermore, we show that the ordered edge state gives rise to an extraordinary surface critical behavior at the (2+1)-dimensional O(3) bulk critical points of the model, which contradicts theoretical predictions based on classical-quantum mapping.
@@ I Introduction
A simple system with SPT order is the spin-1 chain with antiferromagnetic (AF) Heisenberg interaction (often called the Haldane chain), which has a nonzero excitation gap–the Haldane gap, nonlocal string order, and degenerate edge states  [23, 24, 25, 26, 27, 28]. Due to the Haldane gap, the topological nature of the Haldane phase in AF Heisenberg spin-1 chain may remain in (finite) odd-number coupled chains [29] or in quasi-one-dimensional (Q1D) system which is formed by infinite spin-1 chains with weakly but finite inter-chain coupling under certain conditions [30, 31, 32, 33, 34], with gapless edge states along the edge perpendicular to the chain direction [34]. Experimentally, materials described by Q1D spin-1 chains have been found [35, 36].
In many cases, the gapped phase of a spin-1/2 two-leg ladder is directly related to the spin-1 Haldane chain with non-zero string order [37, 28]. Remarkably, a usual ladder with AF rungs ( as shown in Fig. 1 (a)) exhibits the Haldane phase similarly to that of the diagonal ladder, which has AF interchain couplings on plaquette diagonals ( see Fig. 1 (b)), but in different topological sectors.[37]
Inspired by the Q1D weakly coupled spin-1 chains hosting nontrivial SPT order, naturally, one expects that the Q1D weakly coupled spin-1/2 two-leg ladders also present nontrivial SPT states [35, 36]. Despite the similarity to the spin-1-chain Q1D system, the Q1D spin-1/2 ladders hosts a more sophisticated structure in the UV limit, which opens up more possibilities for studying the boundary phenomena that are crucially important for SPT phases. In the previous work [38], we studied the coupled diagonal ladders (CDLs) model and showed that the ground state is indeed an SPT state with gapless edge modes.
@@
The dimer state of the columnar dimerized spin-$1/2$ Heisenberg model on a square lattice (see Fig. 2(a)) is often regarded as a typical example of a topologically trivial state [30, 39]. However, the model can also be viewed as a Q1D coupled usual ladder with AF rungs. Therefore, the gapped dimer state should be an SPT with gapless or degenerate edge modes. In this paper, we will show that this is indeed the case by demonstrating a gapless ferromagnetically (FM) ordered edge state on special zigzag edges, separated by cut1 in Fig. 2(b). On the other hand, the strange correlator has been proposed as a direct probe of the SPT character without cutting the system in Q1D or 2D [40, 33, 41, 42]. We will generalize strange correlators in accordance with the underlying principle of the strange correlator. Utilizing them, we will provide additional evidence that the dimer state is an SPT without cutting the system[40, 33, 41, 42].
@@ II Models and ground state phase diagram
In Ref.[38], we studied the CDLs, which is constructed by coupling the spin-1/2 diagonal ladders [37] to form a 2D lattice, as illustrated in Fig. 2(c) with $J_{\perp}>0$. When interladder coupling is in the region of $0\leq J_{\perp}<(J_{\perp})_{c}=0.17425(3)$, the system stays in the Haldane phase. At $(J_{\perp})_{c}$, the system undergoes a phase transition to the AF phase, which belongs to the (2+1)D O(3) universality class. We found gapless edge modes formed at the ladder endpoints (open circles in Fig. 2(d) ) in the Haldane phase, which is a topologically nontrivial SPT[38].
The ground state phase diagram of the dimerized Heisenberg model, or equivalently, the Q1D coupled usual ladders, has been obtained using the QMC simulations[30], as shown in Fig.3. It is known that the ground state of a usual ladder is a Haldane state, which is a topologically nontrivial SPT state with string order [28]. Similar to the CDLs, the Haldane state is robust against weak higher-dimensional couplings between ladders due to the Haldane gap. The resulting dimer state is also referred to as a Q1D Haldane (Q1DH) state. The transition from the Q1DH state to the AF state belongs to the (2+1)D O(3) universality class. There is no other phase transition when ($J$, $J_{\perp}$) sit on the left of the phase transition line between the AF state and the dimer state, Thus, the Q1DH state is adiabatically connected to the SPT state of the usual ladder. We, therefore, expect the Q1DH state to also be a topologically nontrivial SPT phase. Nontrivial surface/edge states are expected, which may lead to different surface critical behavior compared to the classical phase transition in the 3D O(3) universality class.
@@ III.1 Strange correlator: detect bulk SPT state without a cut
In the simulations of a system of size $L$ with periodic boundaries, we calculate $C_{\rm SC}^{\rm O,E}(L/2)$ by averaging $C_{\rm SC}^{\rm O,E}(i,j)$ at the maximum available distance |ij|=L/2|i-j|=L/2 along an individual ladder, which can be viewed as the strange order parameter for the SPT in the odd and even sectors, respectively.
@@
We then calculate the strange order parameter in 2D. The Q1D coupled ladder systems are strongly anisotropic; we, therefore, set aspect ratio $R=L/L_{\perp}=4$ to reduce corrections to scaling[30] and reach linear size up to $L=352$ along the ladder direction. The CDLs with $J=J_{1}=J_{2}=1$ and weak interladder couplings $J_{\perp}\in[0,0.17425]$ is in the Q1D Haldane phase[38]. We find that the strange order parameter $C_{\rm sc}^{\rm O}(L/2)$ converges to a finite value when $L\to\infty$ in the Q1D Haldane phase, as shown in Fig. 5 (a). This provides direct evidence of a nontrivial SPT order in the Q1D Haldane phase of the weakly coupled diagonal ladders. Similarly, as shown in Fig. 5 (b), the strange order parameter $C_{\rm SC}^{\rm E}(L/2)$ also scales to a finite value when $L\to\infty$ in the dimer phase of the coupled usual ladders, which is also known as the dimersized spin-1/2 Heisenberg model. This result, combined with the fact that the dimer phase is adiabatically connected to the 1D usual ladder, suggests that the dimer phase of the dimerized sipn-1/2 Heisenberg model is an SPT, which is topologically nontrivial.
@@
The hallmark of SPT states is the presence of nontrivial surface states that are either gapless or degenerate in the gapped bulk state[51]. In 1D, this can be understood through the short-range VB picture of the SPT state. As shown in Fig. 1(d) for the diagonal ladder, when the periodic ladder is cut to open two ends, the VB connecting two neighboring $S=1$ spins is broken and leaves two spinons at the ends, leading to four degenerate edge states. For Q1D coupled diagonal ladders, cutting the ladders perpendicular to the ladder direction exposes two boundaries with the VBs connecting neighboring spin-1s cutted but the spin-1s stay untouched. The spinons generated at the ends of the ladders form two dangling spin-1/2 chains at the boundaries, as illustrated in Fig. 2(d), which is gapless. This picture, as expected for an SPT, has been demonstrated numerically in [38].
@@ IV Surface critical behavior
Since the zigzag edge (or surface) is ordered for any $J_{\perp}>0$ due to the bulk-edge correspondence of SPT, which is a quantum mechanical effect, we naturally expect the associated SCB belong to the extraordinary transition, which is characterized by anomalous surface scaling dimension $\eta_{\parallel}$ and the surface magnetic field scaling dimension $y_{h1}$ of the surface field $h_{1}$. In particular, at bulk critical points, the finite-size scaling behaviors of the $C_{\parallel}(L/2)$ and $m_{1}^{2}(L)$ behave as
@@ Acknowledgements.
We acknowledge the helpful discussion with Chao-Ming Jian. Z.W., L.L., and W.G. thank the support from the National Natural Science Foundation of China under Grant No. 12175015. Z.L. and Z.Y. are supported by the Scientific Research Project (No.WU2024B027) and the Start-up Funding of Westlake University. The authors also acknowledge the Super Computing Center of Beijing Normal University and Beijing PARATERA Tech Co., Ltd., and the HPC centre of Westlake University for providing HPC resources. Z.W. is supported by the China Postdoctoral Science Foundation under Grant No.2024M752898. Z.L. thanks the China Postdoctoral Science Foundation under Grants No.2024M762935 and NSFC Special Fund for Theoretical Physics under Grants No.12447119. S.Q.N. is supported by a CRF from the Research Grants Council of the Hong Kong (No. C7037-22GF).

### v2

# Detecting underlying symmetry-protected topological phases via strange correlators and edge engineering

@@ Abstract
The vast majority of symmetry-protected topological (SPT) states are difficult to detect, which often leads to their misidentification as ordinary or topologically trivial phases. In this work, we propose a general framework for detecting these hidden topological states. We distinguish the ordinary matter state from the topological phase by exploiting the boundary effects in space (via surface behaviors on engineered edge) and time (via strange correlators) according to the principle of bulk-edge correspondence. As a concrete example, we study the dimerized spin-1/2 Heisenberg model on a square lattice using quantum Monte Carlo simulations, focusing on its paramagnetic dimer phase and edge states. The dimer phase has been widely regarded as topologically trivial due to its gapped edge state on conventional edges. However, the model can also be viewed as two-dimensional antiferromagnetically (AF) coupled usual ladders, which suggests an SPT state adiabatically connected to the one-dimensional Haldane phase. We resolve this puzzle and demonstrate that the dimer phase is indeed a quasi-one-dimensional SPT state by measuring generalized strange correlators introduced in this work and by showing that the nontrivial gapless edge state on a zigzag edge is ferromagnetically ordered, resulting from effective ferromagnetic interactions between degenerate spinons liberated on each side of the cut. Furthermore, we show that the ordered edge state gives rise to an extraordinary surface critical behavior at the (2+1)-dimensional O(3) bulk critical points of the model, which contradicts theoretical predictions based on classical-quantum mapping. Overall, we establish a standard detection method for uncovering topological phases that masquerade as ordinary states of matter.
@@ I Introduction
A simple system with SPT order is the spin-1 chain with antiferromagnetic (AF) Heisenberg interaction (often called the Haldane chain), which has a nonzero excitation gap–the Haldane gap, nonlocal string order, and degenerate edge states  [23, 24, 25, 26, 27, 28]. Due to the Haldane gap, the topological nature of the Haldane phase in the AF Heisenberg spin-1 chain may remain in (finite) odd-number coupled chains [29] or in quasi-one-dimensional (Q1D) system which is formed by infinite spin-1 chains with weak but finite inter-chain coupling under certain conditions [30, 31, 32, 33, 34], with gapless edge states along the edge perpendicular to the chain direction [34]. Experimentally, materials described by Q1D spin-1 or spin-1/2 chains have been found [35, 36, 37, 38, 39, 40, 40, 41, 42, 43].
In many cases, the gapped phase of a spin-1/2 two-leg ladder [44] is directly related to the spin-1 Haldane chain with non-zero string order [45, 28]. Remarkably, a usual ladder with AF rungs ( as shown in Fig. 1 (a)) exhibits the Haldane phase similarly to that of the diagonal ladder, which has AF interchain couplings on plaquette diagonals ( see Fig. 1 (b)), but in different topological sectors.[45]
Inspired by the Q1D weakly coupled spin-1 chains hosting nontrivial SPT order, naturally, one expects that the Q1D weakly coupled spin-1/2 two-leg ladders also present nontrivial SPT states. Despite the similarity to the spin-1-chain Q1D system, the Q1D spin-1/2 ladders hosts a more sophisticated structure in the UV limit, which opens up more possibilities for studying the boundary phenomena that are crucially important for SPT phases. In the previous works [46, 47], we studied the AF and FM coupled diagonal ladders (CDLs) model, which hosts different SPT phases. The SPT character of the two phases was inferred from their adiabatic connection to the 1D Haldane state and the presence of gapless edge modes.
@@
The columnar dimerized spin-$1/2$ Heisenberg model on a square lattice, realized in Cu compounds such as BaCuSi2O6[41], undergoes a quantum phase transition from the dimer phase to the antiferromagnetic Néel phase (see Fig. 2(a)) and serves as a prototype for quantum criticality[48]. The dimer state has often been regarded as a typical example of a topologically trivial state, as evidenced by its gapped edge states [30, 49]. However, if one views the model as a Q1D coupled usual ladder with AF rungs, then the gapped dimer state could be an SPT, as it is adiabatically connected to the 1D Haldane phase. In this paper, we will solve this puzzle by demonstrating the emergence of a gapless ferromagnetically (FM) ordered edge state on a special zigzag edge, separated by cut1 in Fig. 2(b). Furthermore, the strange correlator has been proposed as a direct probe of the SPT character without cutting the system in Q1D or 2D [50, 33, 51, 52]. To furnish direct evidence of the SPT nature of the dimer state, we will generalize the strange correlator in accordance with its underlying theoretical principle. Utilizing them, we will provide additional evidence that the dimer state is an SPT.
@@ II Models and ground state phase diagram
In Ref.[46, 47], we studied the CDLs, which is constructed by coupling the spin-1/2 diagonal ladders [45] either antiferromagnetically or ferromagnetically to form a 2D lattice, as illustrated in Fig. 2(c) for the AF case $J_{\perp}>0$. When interladder coupling is in the region of $0\leq J_{\perp}<(J_{\perp})_{c}=0.17425(3)$ [46] or $(J_{\perp})_{c}=-0.18271(6)<J_{\perp}<0$ [47], the system stays in the Haldane phases. At $(J_{\perp})_{c}>0$, the system undergoes a phase transition to the AF phase (or to the striped phase at $(J_{\perp})_{c}<0$), which belongs to the (2+1)D O(3) universality class. We found gapless edge modes formed at the ladder endpoints (open circles in Fig. 2(d) ) in the Haldane phases, indicating that the phases are topologically nontrivial SPT [46, 47].
The ground state phase diagram of the dimerized Heisenberg model, or equivalently, the Q1D coupled usual ladders, has been obtained using the QMC simulations[30], as shown in Fig.3. It is known that the ground state of a usual ladder is a Haldane state, which is a topologically nontrivial SPT state with string order [28]. Similar to the CDLs, the Haldane state is robust against weak higher-dimensional couplings between ladders due to the Haldane gap. The resulting dimer state is also referred to as a Q1D Haldane (Q1DH) state. The transition from the Q1DH state to the AF state belongs to the (2+1)D O(3) universality class. There is no other phase transition when ($J$, $J_{\perp}$) sit on the left of the phase transition line between the AF state and the dimer state. Thus, the Q1DH state is adiabatically connected to the SPT state of the usual ladder. We, therefore, expect the Q1DH state to also be a topologically nontrivial SPT phase. Nontrivial surface/edge states are expected, which may lead to different surface critical behavior compared to the classical phase transition in the 3D O(3) universality class.
@@ III.1 Strange correlator: detect bulk SPT state without a cut
In the simulations of a system of size $L$ with periodic boundaries, we calculate $C_{\rm SC}^{\rm O,E}(L/2)$ by averaging $C_{\rm SC}^{\rm O,E}(i,j)$ at the maximum available distance |ij|=L/2|i-j|=L/2 along an individual ladder, which can be viewed as the strange order parameter for the SPT in the odd and even sectors, respectively. (Strictly speaking, the strange correlator is not a genuine order parameter. However, for magnetically disordered states, it serves as an SPT indicator, as discussed in Appendix A).
@@
We then calculate the strange order parameter in 2D. The Q1D coupled ladder systems are strongly anisotropic; therefore, we set the aspect ratio to $R=L/L_{\perp}=4$, where $L=L_{x}$ is the size along the ladder direction and $L_{\perp}=L_{y}$ is the size in the vertical direction. The linear size is reached up to $L=352$. As demonstrated in Appendix B, the critical properties of the AF-dimer transition are not affected by the choice of $R$.
The CDLs with $J=J_{1}=J_{2}=1$ and weak interladder couplings $J_{\perp}\in[0,0.17425]$ is in the Q1D Haldane phase[46, 47]. We find that the strange order parameter $C_{\rm sc}^{\rm O}(L/2)$ converges to a finite value when $L\to\infty$ in the Q1D Haldane phase, as shown in Fig. 5 (a). This provides direct evidence of a nontrivial SPT order in the Q1D Haldane phase of the weakly coupled diagonal ladders. Similarly, as shown in Fig. 5 (b), the strange order parameter $C_{\rm SC}^{\rm E}(L/2)$ also scales to a finite value when $L\to\infty$ in the dimer phase of the coupled usual ladders, which is also known as the dimersized spin-1/2 Heisenberg model. This result, combined with the fact that the dimer phase is adiabatically connected to the 1D usual ladder, suggests that the dimer phase of the dimerized sipn-1/2 Heisenberg model is an SPT, which is topologically nontrivial.
@@
The hallmark of SPT states is the presence of nontrivial surface states that are either gapless or degenerate in the gapped bulk state[60]. In 1D, this can be understood through the short-range VB picture of the SPT state. As shown in Fig. 1(d) for the diagonal ladder, when the periodic ladder is cut to open two ends, the VB connecting two neighboring $S=1$ spins is broken and leaves two spinons at the ends, leading to four degenerate edge states. For Q1D coupled diagonal ladders, cutting the ladders perpendicular to the ladder direction exposes two boundaries with the VBs connecting neighboring spin-1s cutted but the spin-1s stay untouched. The spinons generated at the ends of the ladders form two dangling spin-1/2 chains at the boundaries, as illustrated in Fig. 2(d), which is gapless. This picture, as expected for an SPT, has been demonstrated numerically in [46, 47].
@@ IV Surface critical behavior
Since the zigzag edge (or surface) is ordered for any $J_{\perp}>0$ due to the bulk-edge correspondence of SPT, which is a quantum mechanical effect, we naturally expect the associated SCB to belong to the extraordinary transition, which is characterized by anomalous surface scaling dimension $\eta_{\parallel}$ and the surface magnetic field scaling dimension $y_{h1}$ of the surface field $h_{1}$. In particular, at bulk critical points, the finite-size scaling behaviors of the $C_{\parallel}(L/2)$ and $m_{1}^{2}(L)$ behave as
@@ Acknowledgements.
We acknowledge the helpful discussion with Chao-Ming Jian. Z.W., L.L., and W.G. thank the support from the National Natural Science Foundation of China under Grant Nos. 12175015 and 12574252. Z.L. and Z.Y. are supported by the Scientific Research Project (No.WU2024B027) and the Start-up Funding of Westlake University. The authors also acknowledge the Super Computing Center of Beijing Normal University and Beijing PARATERA Tech Co., Ltd., and the HPC centre of Westlake University for providing HPC resources. Z.W. is supported by the China Postdoctoral Science Foundation under Grant No.2024M752898. Z.L. thanks the China Postdoctoral Science Foundation under Grants No.2024M762935 and NSFC Special Fund for Theoretical Physics under Grants No.12447119. S.Q.N. is supported by a CRF from the Research Grants Council of the Hong Kong (No. C7037-22GF). Y.C.W. acknowledges the support from the Natural Science Foundation of China (Grant No. 12474216) and Zhejiang Provincial Natural Science Foundation of China (Grant No. LZ23A040003), and the Start-up Funding of Hangzhou International Innovation Institute of Beihang University.
Data availability.- The data that support the findings of this article are openly available[65].
Appendix A Strange correlator serves as an SPT indicator only in magnetically disordered phase
For a nontrivial SPT state, the strange correlator either saturates to a constant in 1D or 2D, or at least decays as a power law in 2D[50]. However, this behavior constitutes a necessary but not sufficient condition (an order parameter) for identifying an SPT state.
To illustrate this, we present the strange correlator $C_{\rm SC}^{\rm E}(L/2)$ for the Q1D coupled usual ladders as a function of $J_{1}(=J_{\perp})$ across different system sizes in Fig. 8. At finite $J_{\perp}=J_{1}$, the system is magnetically ordered. We find that within this phase, the strange correlator also converges to a finite value in the thermodynamic limit.
However, as $J_{\perp}=J_{1}\to 0$, the system decouples into topological trivial $S=1/2$ Heisenberg chains, $C_{\rm SC}^{\rm E}(L/2)$ converges to zero. Therefore, for a magnetically disordered phase, the strange correlator can distinguish a quasi-one-dimensional SPT phase and a trivial phase.
Figure 8: The strange correlator $C_{\rm SC}^{\rm E}(L/2)$ plotted as a function of $J_{1}(=J_{\perp})$ for different system sizes in the Q1D coupled usual ladders. The aspect ratio $R=4$, and $J=1$ is fixed.
Appendix B Aspect ratio and critical properties of the dimer-AF transition
The dimerized Heisenberg model is highly anisotropic, where correlation lengths along the ladder direction and the vertical direction may depend on the aspect ratio of the system. We present here additional simulation results of horizontal ($\xi=\xi_{x}$) and vertical ($\xi_{\perp}=\xi_{y}$) correlation lengths in systems with aspect ratios $R=L/L_{\perp}=1$ and $R=4$, respectively, to show that critical properties of the dimer-AF transition are not affected by the choice of aspect ratio. In the simulations, the inverse temperature is set to $\beta=L$, with periodic boundary conditions applied along both $x$ and $y$ lattice directions.
The correlation length is computed from the static spin structure factor (second-moment method [66] ),
$S(\mathbf{q})=\sum_{\mathbf{r}}e^{-i\mathbf{q}\cdot\mathbf{r}}C(\mathbf{r}),$ (11)
where $C(\mathbf{r})=\langle S^{z}_{\mathbf{r}}S^{z}_{\mathbf{0}}\rangle$ is the spin-spin correlation function. The correlation length $\xi=\xi_{x}$ is given by
$\xi=\frac{L}{2\pi}\sqrt{\frac{S(\pi,\pi)}{S(\pi+2\pi/L,\pi)}-1},$ (12)
and $\xi_{\perp}=\xi_{y}$ is given by
$\xi_{\perp}=\frac{L_{\perp}}{2\pi}\sqrt{\frac{S(\pi,\pi)}{S(\pi,\pi+2\pi/L_{\perp})}-1}.$ (13)
Figure 9: Results for $R=1$. (a) $\xi/L$ and (b) $\xi_{\perp}/L_{\perp}$ versus $J_{\perp}$ for different system sizes. Data collapse of $\xi/L$ and $\xi_{\perp}/L_{\perp}$ is shown in (c) and (d), respectively, using the critical exponent $\nu=0.71$ and $J_{\perp c}=0.31407$ for fixed $J=1$. Error bars are much smaller than the symbols.
Any singular physical quantity $A$ (not necessarily divergent) in the vicinity of a QCP obeys the finite-size scaling ansatz
$A(J_{\perp},L)=L^{\kappa}f\left[(J_{\perp}-J_{\perp c})L^{1/\nu},L^{-\omega}\right],$ (14)
where $\nu$ is the correlation length exponent, $f(x)$ is a scaling function, and $\omega>0$ is the effective correction-to-scaling exponent. For $A$ being the scaled correlation length $\xi/L$, or, $\xi_{\perp}/L_{\perp}$, which is dimensionless, we have $\kappa=0$.
Figure 9 (Fig. 10) shows scaled correlation lengths $\xi/L$ and $\xi_{\perp}/L_{\perp}$ of different system sizes as a function of $J_{\perp}$ for $R=1$ and $R=4$, respectively. It is clear that, at the critical point–identified by the crossing of different system sizes–the horizontal and vertical correlation lengths $\xi$ and $\xi_{\perp}$ differ in magnitude for a given aspect ratio $R$. Furthermore, the correlation length $\xi$ and $\xi_{\perp}$ vary with $R$. Nevertheless, the scaled correlation lengths under different settings all collapse onto the scaling function using the universal critical exponent of the (2+1)-dimensional O(3) universality class $\nu=0.71$ [30] and the critical point $J_{\perp c}=0.31407$, indicating that the choice of aspect ratio does not affect the critical properties of the model.
Figure 10: Results for $R=4$. (a) $\xi/L$ and (b) $\xi_{\perp}/L_{\perp}$ versus $J_{\perp}$ for different system sizes $L=L_{x}$. Data collapse of $\xi/L$ and $\xi_{\perp}/L_{\perp}$ is shown in (c) and (d), respectively, using $\nu=0.71$ and $J_{\perp c}=0.31407$ for fixed $J=1$. Error bars are much smaller than the symbols.